### v1

@@ 4 Methods
Our knowledge transfer framework is based on the following concept (Fig. 1): Any contribution by a participant of the network is associated with a task fingerprint, i.e., a shareable task representation that respects individual privacy regulation (e.g., that the data itself must not be shared). Meta information about suitable training design decisions, such as model architecture or augmentation policy, is stored along with the fingerprint. Optionally, the data may be tagged as shareable. To learn from the current network of tasks, a user generates the fingerprint for their task and queries the knowledge cloud. Based on the most relevant tasks in the pool according to fingerprint matching, relevant training strategies and data can be retrieved. The retrieved meta information and data are used to compile a training pipeline with different components of transferred knowledge. This section details our core methodological contributions as well as the evaluation concept we developed to assess the value of our approach.
@@ 4.1 Task fingerprinting
where $\operatorname{soft}$ is the softmax function such that $\hat{q}_{i}^{(j)}=\exp q_{i}^{(j)}/\sum_{j}\exp q_{i}^{(j)}$. The softmax is used to circumvent potentially empty bins. The key novelty of our approach is the combination of a sufficiently granular task embedding (compared to P2L [25]) with a flexible weighting to guide feature attention (compared to VDNA [17]). Based on our exploratory experiments on the development tasks (see Fig. 3), we identified three favorable settings of hyperparameters: bKLD(small,target) a ”small” fingerprint with $b=100$ bins and weighting by the softened feature average of the target task $w=\operatorname{soft}(\sum_{i}^{n}\mathbf{v}_{i}/n)$, bKLD(large,source) with $b=1,000$ bins and weighting by normalized feature average of the source task as well as bKLD(large,unweighted) with $b=1,000$ and uniform weighting. To compare bKLD with the manual task matching approach, we extracted keywords from the semantic description of tasks, e.g., imaging modality, anatomical regions, and entities of interest, and computed the intersection over union between any pair of tasks. Ties were resolved by preferring larger source tasks.
@@ 4.3 Validation metrics
Improvement measures the difference in applying the knowledge transfer via $s$ compared to the baseline: $m(t_{\mathfrak{S}}(\mathfrak{T}))-m(t(\mathfrak{T}))$. Percentile measures the relative position of $m(t_{\mathfrak{S}}(\mathfrak{T}))$ within all potential transfer outcomes: |$|\{o\in\mathbf{O}:o\leq m(t_{\mathfrak{S}}(\mathfrak{T}))\}|/|\mathbf{O}|$. Regret [49] measures how much of the remaining performance gap (from the performance determined by $s$ to optimal performance) could be overcome by making the optimal choice within $\mathbf{P}$: $[\max\mathbf{O}-m(t_{\mathfrak{S}}(\mathfrak{T}))]/[1-m(t_{\mathfrak{S}}(\mathfrak{T}))]$. Gain [20] is a weaker form of improvement and returns 0 for a negative transfer, otherwise 1.
@@ 4.5.1 Single task baseline
The single task baseline comprises a “transfer-free” training recipe and subsequently simulates a model performance that is indicative of an achievable result without significant resource investment in experimentation. Coming up with such a fair and meaningful training scheme for each task presents a major challenge given the large heterogeneity between them. Too much individual optimization per task is costly and introduces another potential factor of bias; on the other hand, applying a uniform training scheme on all tasks might lead to unrealistic performances. We aimed for a balance with the following strategy: For faster convergence and more stable training, we use ImageNet [2] pretrained models throughout [6]. Starting from the recommendations given by Wightman et al. [86], we made slight adaptations to the training pipeline to ensure model convergence, avoid overfitting, and generally improve performance on $\mathbf{P}_{\text{dev}}$. On an individual task level, we used the automated tuning of the learning rate [87] that is implemented in Pytorch Lightning [88]. As soon as we identified a solid configuration, we kept it for all ensuing experiments. The unmodified pipeline served as a baseline to compute gain and improvement. To compute the other meta metrics, it was necessary to train all potential knowledge transfer-induced pipeline adaptations (on all tasks) - this allowed us to rank all source tasks (percentile and weightedtau) and deduce the theoretically optimal performance (regret).
@@ Acknowledgements
The authors would like to thank Maike Rees for her help on designing Fig. 1 and Fig. 3. We would also like to thank all members of the ”Intelligent Systems in Surgical Endoscopy” team for fruitful discussions. We are grateful for the linguistic corrections by Minu Dietlinde Tizabi and Marcel Knopp. This project has been funded by the German Federal Ministry of Health under the reference number 2520DAT0P1 as part of the pAItient (Protected Artificial Intelligence Innovation Environment for Patient Oriented Digital Health Solutions for developing, testing and evidence based evaluation of clinical value) project. The funder played no role in study design, data collection, analysis and interpretation of data, or the writing of this manuscript. The present contribution is supported by the Helmholtz Association under the joint research school “HIDSS4Health – Helmholtz Information and Data Science School for Health”.
@@
P.G. and L.M.H. conceptualized the work. P.G., A.S., T.A. and L.M.H. designed the study. P.G. and A.S. performed the coding and data preparation. P.G. trained all deep learning models and conducted the statistical analysis. T.A. and L.M.H. provided advisory support for the project. P.G. and L.M.H. prepared the initial draft of the manuscript, with all authors participating in the refinement.

### v2

@@ 4 Methods
Our knowledge transfer framework is based on the following concept (Fig. 1): Any contribution by a participant of the network is associated with a task fingerprint, i.e., a shareable task representation that respects individual privacy regulation (e.g., that the data itself must not be shared). Meta information about suitable training design decisions, such as model architecture (e.g., a timm [61] string) or augmentation policy (e.g., an albumentations [62] serialization), is stored alongside the fingerprint. Optionally, the data may be tagged as shareable. To learn from the current network of tasks, a user generates the fingerprint for their task and queries the knowledge cloud. Based on the most relevant tasks in the pool according to fingerprint matching, relevant training strategies and data can be retrieved. The retrieved meta information and data are used to compile a training pipeline with different components of transferred knowledge. This section details our core methodological contributions as well as the evaluation concept we developed to assess the value of our approach.
@@ 4.1 Task fingerprinting
where $\operatorname{soft}$ is the softmax function such that $\hat{q}_{i}^{(j)}=\exp q_{i}^{(j)}/\sum_{j}\exp q_{i}^{(j)}$. The softmax is used to circumvent potentially empty bins. The key novelty of our approach is the combination of a sufficiently granular task embedding (compared to P2L [34]) with a flexible weighting to guide feature attention (compared to VDNA [20]).
4.2 Hyperparameter optimization
During the development phase, we began by exploring variations of previously proposed methods for task similarity [20, 15, 33, 34, 35, 65, 31]. We combined various embedding mappings $f$, including the binning scheme, with multiple distance measures $d$. We examined several smoothing functions (e.g., normalization, softmax, and symmetric uniform smoothing) for both the source and target embeddings, as well as weighting schemes. The list of all configurations can be found in our code repository. Based on our exploratory experiments on the development tasks (see Fig. 3), we identified two favorable settings of hyperparameters: bKLD(small) a ”small” fingerprint with $b=100$ bins and weighting by the softened feature average of the target task $w=\operatorname{soft}(\sum_{i}^{n}\mathbf{v}_{i}/n)$ as well as bKLD(large with $b=1,000$ and uniform weighting. To compare bKLD with the manual task matching approach, we extracted keywords from the semantic description of tasks, e.g., imaging modality, anatomical regions, and entities of interest, and computed the intersection over union between any pair of tasks. Ties were resolved by preferring larger source tasks.
@@ 4.4 Validation metrics
Improvement measures the difference in applying the knowledge transfer via $s$ compared to the baseline: $m(t_{\mathfrak{S}}(\mathfrak{T}))-m(t(\mathfrak{T}))$. Percentile measures the relative position of $m(t_{\mathfrak{S}}(\mathfrak{T}))$ within all potential transfer outcomes: $|\{o\in\mathbf{O}:o\leq m(t_{\mathfrak{S}}(\mathfrak{T}))\}|/|\mathbf{O}|$. Regret [67] measures how much of the remaining performance gap (from the performance determined by $s$ to optimal performance) could be overcome by making the optimal choice within $\mathbf{P}$: $\max\mathbf{O}-m(t_{\mathfrak{S}}(\mathfrak{T}))]/[1-m(t_{\mathfrak{S}}(\mathfrak{T}))]$. Gain [28] is a weaker form of improvement and returns 0 for a negative transfer, otherwise 1.
@@ 4.6 Validation of knowledge transfer quality
To facilitate comparison, we included a selection of previously proposed task selectors based on task distances as baselines. These were chosen to balance “classic” methods, such as FID [32, 33]), with more recent methods, such as VDNA [20]). We also compared bKLD with approaches that use more dense task representations, such as Sinkhorn Divergence [35], as well as approaches that use much simpler representations, such as P2L [34]). FED [31, 15] complements these baselines because it is not based on feature extraction. The “manual” baseline was intended to capture a common and tangible approach to the problem.
@@ 4.6.1 Single task baseline
The single task baseline comprises a “transfer-free” training recipe and subsequently simulates a model performance that is indicative of an achievable result without significant resource investment in experimentation. Coming up with such a fair and meaningful training scheme for each task presents a major challenge given the large heterogeneity between them. Too much individual optimization per task is costly and introduces another potential factor of bias; on the other hand, applying a uniform training scheme on all tasks might lead to unrealistic performances. We aimed for a balance with the following strategy: For faster convergence and more stable training, we use ImageNet [2] pretrained models throughout [6]. Starting from the recommendations given by Wightman et al. [104], we made slight adaptations to the training pipeline to ensure model convergence, avoid overfitting, and generally improve performance on $\mathbf{P}_{\text{dev}}$. On an individual task level, we used the automated tuning of the learning rate [105] that is implemented in Pytorch Lightning [106]. Medical data often suffers from class imbalance, which can affect model performance. To address this important issue, we worked with a variety of imbalanced tasks, measuring imbalance as the ratio of the size of the largest class to the size of the smallest class, with ratios up to 171. Class-balanced sampling was used to show the models the less frequent classes with a higher probability. The pipeline was optimized collectively across tasks during development. Afterwards, it was validated on the validation tasks. It is worth noting that this approach occasionally results in non-convergence during training and frequently performs less effectively than a training procedure that incorporates more task-specific pipeline tuning. However, we argue that the primary purpose of task fingerprinting is not to immediately produce a state-of-the-art model, but rather to refine the model development process by utilizing advantageous initialization and assisted iteration. Second, in order to validate the methodology, it is sufficient to demonstrate positive trends in a suboptimal yet sensible regime. As soon as we identified a solid configuration, we kept it for all ensuing experiments. The unmodified pipeline served as a baseline to compute gain and improvement. To compute the other meta metrics, it was necessary to train all potential knowledge transfer-induced pipeline adaptations (on all tasks) - this allowed us to rank all source tasks (percentile and weightedtau) and deduce the theoretically optimal performance (regret).
@@ 4.6.5 Scenario 4: Co-Training Data
4.6.6 Generalization Scenario
To investigate the generalization of our methodology to different modalities and tasks prospectively, we applied the concept developed on 2D classification tasks to 3D segmentation tasks. To this end, we leveraged one of the most widely used datasets, namely the Medical Segmentation Decathlon challenge data [42], and official results [41]. For feature extraction we used a Masked Autoencoder (MAE) [40] pretrained by Eckstein et al. [130] on two large 3D datasets [131, 132] based on the implementation by Wald et al. [133] and a ResEncL backbone [134]. Again, we extracted 10,000 feature maps per dataset from the bottleneck layer of the network, without masking the input patch. The feature maps originally had a size of $320\times 5\times 5\times 5$, however for computational efficiency, we reduced all but the first dimension via the arithmetic mean. The patches were sampled randomly across patients, available imaging modalities, and spatial locations. The challenge provides the validation values for 19 participating teams, on 17 tasks for two performance measures: Dice Similarity Coefficient (DSC) and Normalized Surface Dice (NSD). When analyzing the results, however, we observed that the distribution of ”best models per task” was skewed (a single participant won 29 out of the 34 rankings). Hence, applying our methodology for scenario 1, which only considers the best performance on the source task, would lead to a trivial setup, where almost all source tasks would be optimal for most target tasks. To address this issue, we computed the transferability between tasks as the Pearson correlation coefficient of the participants’ results on two target tasks. Note that we excluded two participants who occasionally performed with a value of zero, indicating algorithmic failure.
@@ Acknowledgements
The authors would like to thank Maike Rees for her help on designing Fig. 1 and Fig. 3. We would also like to thank all members of the ”Intelligent Systems in Surgical Endoscopy” team for fruitful discussions. We are grateful for the linguistic corrections by Minu Dietlinde Tizabi and Marcel Knopp. This project has been funded by the German Federal Ministry of Health under the reference number 2520DAT0P1 as part of the pAItient (Protected Artificial Intelligence Innovation Environment for Patient Oriented Digital Health Solutions for developing, testing and evidence based evaluation of clinical value) project. The funder played no role in study design, data collection, analysis and interpretation of data, or the writing of this manuscript. The present contribution is supported by the Helmholtz Association under the joint research school “HIDSS4Health – Helmholtz Information and Data Science School for Health”. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (project NEURAL SPICING grant agreement No. 101002198).
@@
P.G. and L.M.H. conceptualized the work. P.G., A.S.,C.U., T.A., K.M.H., and L.M.H. designed the study. P.G., C.U., and A.S. performed the coding and data preparation. P.G. trained all deep learning models and conducted the statistical analysis. T.A., K.M.H., and L.M.H. provided advisory support for the project. P.G. and L.M.H. prepared the initial draft of the manuscript, with all authors participating in the refinement.